y=x+1
-x+y=3
solve the linear equations

Answer by richwmiller(17219)

y=x+1
-x+y=3
substitute x+1 for y
-x+(x+1)=3
1=3
These equations have no solution
rearrange the second
y=x+3
y=x+1
They are parallel with the same slope.